OpenStudy (anonymous):

last question 4/10+?/5=3/5

OpenStudy (anonymous):

what is ?

myininaya (myininaya):

4/10=2/?

myininaya (myininaya):

4/10=2/5 right?

OpenStudy (anonymous):

i dont no i need to find that out i thought maybe 2

myininaya (myininaya):

4/10 4 and 10 are both even so that means they are both divisible by 2 4 divided by 2=2 10 divided by 2=5 so 4/10=2/5 so what is the question mark in: 2/5+?/5=3/5

OpenStudy (anonymous):

ok then i had it right not sure about this stuff thanks

myininaya (myininaya):

since the denominators are the same on both the left hand side and the right hand side then the sum of the numerators of the left hand side equals the numerator of the right hand side so the 2/5+1/5=3/5